use the quadratic formula to solve the equation to the exact value or round to two decimal places. x2+x=5
Answer by rothauserc(4718)

x^2 +x -5 = 0
:
the standard form of a quadratic equation is
:
ax^2 +bx +c = 0
:
for this problem a=1, b=1, c=-5
:
the quadratic formula is
:
x = (-b + or - sqrt(b^2-4ac))/2a
:
x = (-1 +sqrt(1^2 -4(1)(-5))/2(1) = 1.7912 approximately 1.79
:
x = (-1 -sqrt(1^2 -4(1)(-5))/2(1) = -2.7913 approximately -2.79
: